# Nonlocal pseudo-symmetries and Bäcklund transformations as $\mathcal{C}$-morphisms

### v2

@@
Nonlocal pseudo-symmetries and Bäcklund transformations as $\mathcal{C}$-morphisms
Diego Catalano Ferraioli1,  Tarcísio Castro Silva2
?abstractname?
In this paper, we show how factorization with respect to nonlocal pseudosymmetries allows one to obtain Bäcklund transformations, interpreted as nonlocal $\mathcal{C}$-morphisms of differential equations. This approach is illustrated through several examples.
@@ 1 Introduction
In its modern formulation, an auto-Bäcklund transformation for a differential equation $\mathcal{E}$ refers to a map that associates, to any given solution of $\mathcal{E}$, a family of new solutions of $\mathcal{E}$, typically by integrating a system of ODEs. Then, more generally, a Bäcklund transformation is understood as a transformation that, via the integration of a system of ordinary differential equations, maps solutions of one differential equation $\mathcal{E}$ to solutions of another equation $\mathcal{Y}$.
@@
Given a nonlinear differential equation, the problem of determining the existence of its auto-Bäcklund transformations—or, more generally, of Bäcklund transformations between this equation and other equations of interest—is a significant problem that has attracted the attention of many mathematicians (see, for instance [4, 24, 25, 29, 33, 35, 37, 45] and references therein). However, it is a difficult problem which is still open in its full generality, despite having been studied for a long time and, in recent years, interesting progress having been made in some particular cases (see, for instance, the works [2, 3, 16, 17, 18, 42, 43]).
@@
On the other hand, a notion closely related to that of a Bäcklund transformation is that of a differential substitution. In fact, a differential substitution from an equation $\mathcal{E}$ to an equation $\mathcal{Y}$ is a transformation that maps solutions of $\mathcal{E}$ to solutions of $\mathcal{Y}$, without requiring the integration of a system of ordinary differential equations. Classical examples of differential substitutions include the Cole–Hopf transformation, which maps solutions of the heat equation to solutions of Burgers equation, and the Miura transformation, which maps solutions of the modified Korteweg-de Vries (mKdV) equation to solutions of the Korteweg-de Vries (KdV) equation.
@@
As with Bäcklund transformations, given a differential equation $\mathcal{E}$, it is often of great interest to know whether a differential substitution exists that maps its solutions into those of another equation $\mathcal{Y}$. Regarding this issue, in the context of evolutionary equations, Sokolov took a significant step forward in [38], highlighting a fundamental relationship between factorization via pseudosymmetries and differential substitutions. However, although Sokolov underscored the importance of pseudosymmetries through this discovery, these have been surprisingly underexplored in the literature. Indeed, aside from some further contributions by Sokolov himself, to the best of our knowledge this notion has been explicitly addressed only in few papers (e.g., [14, 15, 29]), and, in simplified forms, in studies concerning mainly the reduction of ordinary differential equations (see, for instance, [10, 11, 12, 32] and the references therein).
@@
As in the case of common symmetries, we can refer to pseudosymmetries of a covering $\tilde{\mathcal{E}}$ of an equation $\mathcal{E}$ as nonlocal pseudosymmetries of $\mathcal{E}$. This further type of extension of the notion of pseudosymmetry plays a fundamental role in this paper, together with the notion of $\mathcal{C}$-morphism and the factorization of an equation by means of pseudosymmetries. Indeed, in the present paper we show how the factorization by nonlocal pseudosymmetries allows us to obtain Bäcklund transformations between differential equations.
@@
The paper is organized as follows. Section 2 is a section of preliminaries, devoted to some general aspects of the geometry of differential equations, viewed as submanifolds of jet spaces, and a discussion of more specific notions that will play a particularly significant role in the remainder of the paper. This section is divided into the following four subsections: the first is devoted to a brief introduction to differential equations as submanifolds of jet spaces; the second deals with $\mathcal{C}$-morphisms; the third, after a short review of the notion of symmetry, provides a discussion of the notion of pseudosymmetry; and finally the fourth subsection reviews the theory of differential coverings. In particular, while discussing the notion of pseudosymmetry, a description of the factorization process through pseudosymmetries is also provided. Furthermore, some significant examples are included in this section to illustrate the central notions discussed here. Next, in Section 3, we demonstrate how starting from zero-curvature representations of an equation $\mathcal{E}$, we can determine Riccati-type differential coverings for $\mathcal{E}$, together with nonlocal conservation laws. This type of covering, with the related conservation laws, is particularly useful for determining possible nonlocal pseudosymmetries of $\mathcal{E}$. Finally, Section 4 shows how factorization using nonlocal pseudosymmetries of an equation $\mathcal{E}$ can provide Bäcklund transformations for $\mathcal{E}$. This approach to determining Bäcklund transformations is illustrated here through a series of examples.
@@
manifolds and maps are assumed to be of $C^{\infty}$ class;
@@
for jet bundles $J^{k}(\pi)$ and $J^{\infty}(\pi)$ of a fiber bundle $\pi:E\rightarrow M$ (in particular for total derivatives, Cartan distributions, prolongations, symmetries, horizontal and vertical differentials, conservation laws, zero-curvature representations, and differential coverings) we will use the notations and conventions revised in Subsections2.1, 2.3.1 and 2.4, which follow references like [8, 22, 28, 34, 40];
@@
by considering a system of differential equations $\mathcal{E}\subset J^{k}(\pi)$, we often describe it as =$\mathcal{E}=\{F^{1}(x,u_{\sigma})=0,...,F^{r}(x,u_{\sigma})=0\}$ except when a differential covering of $\mathcal{E}$ is involved in the discussion;
when $\mathcal{E},\mathcal{Y}\subset J^{k}(\pi)$ are such that $\mathcal{Y}$ defines a $q$-dimensional differentiable covering (or extension) of $\mathcal{E}$, we will usually distinguish the fiber coordinates $\{u^{1},...,u^{m}\}$ into the two subsets $\{z^{1},...,z^{m-q}\}$ (also referred to as local variables) and $\{v^{1},...,v^{q}\}$ (also referred to as nonlocal variables), such that =$\mathcal{E}=\{F^{1}(x,z_{\sigma})=0,...,F^{r}(x,z_{\sigma})=0\}$ and $\mathcal{Y}=\{F^{h}(x,z_{\sigma})=0,\,v_{i}^{s}=X_{i}^{s}(x,z_{\sigma},v)\}$, where $h=1,...,r$, $s=1,...q$ and $i$ enumerates the independent variables $x_{1},...,x_{n}$;
for differentiable covering of Riccati type, we will usually denote nonlocal variables by $\{\rho^{j}\}$;
by a “basic system of invariants” we mean a system of invariants that generates all invariants through prolongations and functional combinations, while by a “complete system of invariants” we mean a system of invariants that generates a particular class of invariants (e.g., those of a specific order) through functional combinations.
@@ 2 Preliminaries
We assume the reader is familiar with the geometric theory of differential equations. However, to make the main results accessible to a wide range of readers, we collected some notations and basic facts of this theory used throughout the paper. We also introduce lesser-known topics that will be key in the forthcoming sections. The reader is referred to [8, 22, 34] and [4, 20, 38, 40], and references therein, for further details.
@@ 2.1 Differential equations as submanifolds of jet spaces
Let $\pi:E\rightarrow M$ be a fiber bundle, with $dim\,M=n$ and $dim\,E=n+m$. For any $k\in\mathbb{N}$ we denote by $J^{k}(\pi)$ the manifold of $k$-th order jets of sections of $\pi$ and by $\pi_{k}:J^{k}(\pi)\rightarrow M$ the $k$-order jet bundle of sections of $\pi$. By definition, a point $\theta$ of $J^{k}(\pi)$ is an equivalence class $\theta=[s]_{a}^{k}$ of smooth sections $s$ of $\pi$ whose graphs at $a\in M$ pass through the same point $s(a)$, where they have the same contact up to order $k$. Hence, any section $s$ of $\pi$ togheter with its derivatives up to order $k$, determines a section $j_{k}(s)$ of $\pi_{k}$ which is called the $k$-th order prolongation of $s$. For $h>k$ we denote by $\pi_{h,k}:J^{h}(\pi)\rightarrow J^{k}(\pi)$ the natural projection of $J^{h}(\pi)$ onto $J^{k}(\pi)$, given by $[s]_{a}^{h}\rightarrow[s]_{a}^{k}$.

### v5

# Nonlocal pseudosymmetries and Bäcklund transformations as $\mathcal{C}$-morphisms

@@
Abstract
In this paper, we show how factorisation with respect to nonlocal pseudosymmetries allows one to obtain Bäcklund transformations, interpreted as nonlocal $\mathcal{C}$-morphisms of differential equations. According to this approach, which is illustrated through several examples, the Bäcklund transformations are determined by basic invariants of the exploited nonlocal pseudosymmetries.
@@ 1 Introduction
In its modern formulation, an auto-Bäcklund transformation for a differential equation $\mathcal{E}$ refers to a map that associates, to any given solution of $\mathcal{E}$, a family of new solutions of $\mathcal{E}$, typically by integrating a system of ODEs. Then, more generally, a Bäcklund transformation is understood as a transformation that, via the integration of a system of ordinary differential equations, maps solutions of one differential equation $\mathcal{E}$ to solutions of another equation $\mathcal{E}^{\prime}$.
@@
Given a nonlinear differential equation, the problem of determining the existence of its auto-Bäcklund transformations—or, more generally, of Bäcklund transformations between this equation and other equations of interest—is an important problem that has attracted the attention of many mathematicians (see, for instance [4, 27, 28, 36, 38, 40, 48] and references therein). However, it is a difficult problem which is still open in its full generality, despite having been studied for a long time and, in recent years, interesting progress having been made in some particular cases (see, for instance, the works [2, 3, 18, 19, 20, 45, 46]).
@@
On the other hand, a notion closely related to that of a Bäcklund transformation is that of a differential substitution. In fact, a differential substitution from an equation $\mathcal{E}$ to an equation $\mathcal{E}^{\prime}$ is a transformation that maps solutions of $\mathcal{E}$ to solutions of $\mathcal{E}^{\prime}$, without requiring the integration of a system of ordinary differential equations. Classical examples of differential substitutions include the Cole–Hopf transformation, which maps solutions of the heat equation to solutions of Burgers equation, and the Miura transformation, which maps solutions of the modified Korteweg-de Vries (mKdV) equation to solutions of the Korteweg-de Vries (KdV) equation.
@@
As with Bäcklund transformations, given a differential equation $\mathcal{E}$, it is often of great interest to know whether a differential substitution exists that maps its solutions into those of another equation $\mathcal{E}^{\prime}$. Regarding this issue, in the context of evolutionary equations, Sokolov took a significant step forward in [41], highlighting a particularly important relationship between factorisation via pseudosymmetries and differential substitutions. However, although Sokolov underscored the importance of pseudosymmetries through this discovery, these have been surprisingly underexplored in the literature. Indeed, aside from the works of Sokolov himself, pseudosymmetries have been little studied in their full generality. In a simplified form, they have been used in several works mainly to reduce and integrate ordinary differential equations (see, for example, [10, 11, 13, 35] and the references therein). While for the particular case of integrable pseudosymmetries, some possible uses have been discussed in the study of differential coverings [16, 17] (see [26] for the notion of differentiable covering) or as a tool to explore the effectiveness and generating power of an already known Bäcklund transformation [32]. Therefore, no previous work on using pseudosymmetries addresses the problem of determining Bäcklund transformations.
@@
As in the case of common symmetries, we can refer to pseudosymmetries of a covering $\tilde{\mathcal{E}}$ of an equation $\mathcal{E}$ as nonlocal pseudosymmetries of $\mathcal{E}$. This type of extension of the notion of pseudosymmetry plays a fundamental role in this paper, together with the notion of $\mathcal{C}$-morphism and the factorisation of an equation by means of pseudosymmetries. Indeed, in the present paper, we show how the factorisation by nonlocal pseudosymmetries allows one to obtain Bäcklund transformations, understood as $\mathcal{C}$-morphisms determined by basic invariants of the exploited pseudosymmetries. In particular, following this approach, for equations admitting zero-curvature representations (ZCRs), it is natural to consider the nonlocal pseudosymmetries defined in a kind of Riccati-type differentiable coverings, which are shown to be determined by such representations. Indeed, using this kind of covering, various examples of Bäcklund transformations, for equations admitting ZCRs, can all be obtained similarly, in a unified way, by factorisation with respect to nonlocal pseudosymmetries.
@@
The paper is organised as follows. Section 2 is a section of preliminaries, devoted to some general aspects of the geometry of differential equations and a discussion of more specific notions that will play a particularly significant role in the remainder of the paper. This section is divided into the following four subsections: the first is devoted to a brief introduction to differential equations as submanifolds of jet spaces; the second deals with $\mathcal{C}$-morphisms; the third, after a short review of the notion of symmetry, provides a discussion of the notion of pseudosymmetry together with a brief description of how this notion leads to a factorisation procedure; finally the fourth subsection reviews the theory of differential coverings. In particular, some significant examples are included in this section to illustrate the central notions discussed here. Next, starting from ZCRs of $\mathcal{E}$, in Section 3 we show how to determine Riccati-type differential coverings of $\mathcal{E}$ and corresponding nonlocal conservation laws. This type of covering, with the related conservation laws, is particularly useful for determining possible nonlocal pseudosymmetries of $\mathcal{E}$. Finally, Section 4 shows how factorisation using nonlocal pseudosymmetries of an equation $\mathcal{E}$ can provide Bäcklund transformations for $\mathcal{E}$. This approach to determining Bäcklund transformations is illustrated here through a series of examples.
@@
all objects in the paper, e.g., manifolds, mappings, functions, vector fields , etc, are supposed to be smooth;
@@
for jet bundles $J^{k}(\pi)$ and $J^{\infty}(\pi)$ of a fiber bundle $\pi:E\rightarrow M$ (in particular for total derivatives, Cartan distributions, prolongations, symmetries, horizontal and vertical differentials, conservation laws, zero-curvature representations, and differential coverings) we will use the notations and conventions revised in Subsections 2.1, 2.3.1 and 2.4;
@@
the algebra of smooth functions on a manifold $N$ is usually denoted by $C^{\infty}(N)$, whereas $\mathcal{F}_{k}(\pi)$ and $\mathcal{F}(\pi)$ will denote the algebra of smooth functions on $J^{k}(\pi)$ and $J^{\infty}(\pi)$, respectively;
by considering a system of differential equations $\mathcal{E}\subset J^{k}(\pi)$, we usually describe it as $\mathcal{E}=\{F^{1}(x,u_{\sigma})=0,...,F^{r}(x,u_{\sigma})=0\}$;
when $\mathcal{E},\mathcal{V}\subset J^{k}(\pi)$ are such that $\mathcal{V}$ is a $q$-dimensional differentiable extension of $\mathcal{E}$, i.e., $\mathcal{V}$ defines a $q$-dimensional differentiable covering $\mathcal{V}^{(\infty)}\rightarrow\mathcal{E}^{(\infty)}$, we will usually distinguish the fiber coordinates $\{u^{1},...,u^{m}\}$ of $\pi$ into the two subsets $\{z^{1},...,z^{m-q}\}$ (also referred to as local variables) and $\{v^{1},...,v^{q}\}$ (also referred to as nonlocal variables), such that $\mathcal{E}=\{F^{1}(x,z_{\sigma})=0,...,F^{r}(x,z_{\sigma})=0\}$ and $\mathcal{V}=\{F^{h}(x,z_{\sigma})=0,\,v_{i}^{s}=X_{i}^{s}(x,z_{\sigma},v)\}$, where h=1,,rh=1,...,r, s=1,qs=1,...q and $i$ enumerates the independent variables $x_{1},...,x_{n}$;
for Riccati-type differentiable coverings of an equation $\mathcal{E}$, we will usually denote nonlocal variables by $\{\rho^{j}\}$ and the differentiable extension of $\mathcal{E}$ by $\mathcal{Y}$;
by a “basic system of invariants”, we mean a system of invariants that generates all invariants through prolongations and functional combinations, whereas by a “complete system of invariants”, we mean a system of invariants that generates all invariants only through functional combinations.
@@ 2 Preliminaries
We assume the reader is familiar with the geometric theory of differential equations. However, to make the main results accessible to a wide range of readers, we collected here some notations and basic facts of this theory used throughout the paper. We also introduce lesser-known topics that will be key in the forthcoming sections. The reader is referred to [8, 25, 37, 43] and [4, 12, 21, 23, 31, 41], and references therein, for further details.
@@ 2.1 Differential equations as submanifolds of jet spaces
Consider a fiber bundle $\pi:E\rightarrow M$, with dimM=ndim\,M=n and dimE=n+mdim\,E=n+m. Given a smooth (local) section $s$ of $\pi$ at $a\in M$, for any $k\in\mathbb{N}$, we denote by $s]_{a}^{k}$ the $k$-th order jet of $s$ at $a$; by definition $s]_{a}^{k}$ is the equivalence class of (local) smooth sections of $\pi$ that are $k$-order tangent (or, have $k$-fold contact) to $s$ at $a$. The space $J^{k}(\pi)$ of $k$-th order jets of sections of $\pi$ is naturally equipped with a differentiable manifold structure, induced by bundle atlas of $\pi$. If $\{x_{1},...,x_{n}\}$ are local coordinates on $M$ and $\{u^{1},...,u^{m}\}$ local fiber coordinates of $\pi$, the induced canonical coordinates on $J^{k}(\pi)$ will be denoted by $\{x_{i},u_{\sigma}^{j}\}$, where $i\in\{1,...,n\}$, $j\in\{1,...,m\}$ and $\sigma=(\sigma_{1},...,\sigma_{n})$ is a multi-index of order $|\sigma|=\sigma_{1}+...+\sigma_{n}$ such that $0\leq|\sigma|\leq k$; by definition, if $\theta=[s]_{a}^{k}$, then $x_{i}(\theta):=x_{i}(a)$ and $u_{\sigma}^{j}(\theta):=\frac{\partial^{|\sigma|}s^{j}}{\partial x_{1}^{\sigma_{1}}...\partial x_{n}^{\sigma_{n}}}(a)$. The manifold $J^{k}(\pi)$ is usually referred to as the $k$-order jet bundle of sections of $\pi$, since for any $k\in\mathbb{N}$ the natural projection $\pi_{k}:J^{k}(\pi)\rightarrow M$, $s]_{a}^{k}\mapsto a$, is a fiber bundle; in canonical coordinates, $\pi_{k}$ has the form $(x_{i},u_{\sigma}^{j})\mapsto(x_{i})$. In particular, $J^{0}(\pi)$ can be identified with $E$ and $\pi_{0}$ with $\pi$.